The perimeter for a sports field is 292ft.the length is 48ft longer than the width.
what is the width of the field? what is the length of the field?

Answer by Alan3354(69443)

Do it like this one:
----------------------
The length of a rectangle is 10 cm longer than its width. The perimeter of the rectangle is 68 cm. What are the length and width of the rectangle?
-------
Perimeter = 2W + 2L = 68
W + L = 34
L = W + 10
Sub for L in the 1st eqn
W + (W+10) = 34
2W + 10 = 34
W = 12 cm